What force is necessary to keep a body of mass 0.8kg moving in a horizontal circle of radius 0.7m with period 0.5s.


1
Expert's answer
2021-08-06T15:26:12-0400

Given:

"m=0.8\:\rm kg"

"r=0.7\:\rm m"

"T=0.5\:\rm s"

The Newton's second law says

"F=ma=\frac{mv^2}{r}=\frac{4\pi^2rm}{T^2}"

"F=\frac{4\pi^2*0.7*0.8}{0.5^2}=88\:\rm N"